# Minimum Cost Loop Nests for Contraction of a Sparse Tensor with a Tensor Network

### v1

@@ 4.2.1. Formal Definition of Contraction Path and Loop Nest
For a contraction of $N+1$ tensors, a contraction path is given by a depth-first postordering of a $12N+$ 1-node binary tree $T$ where the $N$ leaves are the input tensors, and each internal node corresponds to the contraction of a pair of input tensors and/or intermediates, so all non-leaf nodes have exactly two children. We represent a contraction path by the tree $T$ and an ordered collection of index set 3-tuples, $L=(L_{1},\dots,L_{N})$, where each $L_{i}$ contains the indices of the tensor operands and output at each of the $N$ internal tree nodes.
Note that while a contraction path is defined above based on a tree, this tree is different tree from the one we seek to construct a loop nest. In a loop nest tree, each node corresponds to a loop and each leaf is a term in the contraction path. We represent this tree implicitly by leveraging the one-to-one correspondence between a fully-fused loop nest tree and a choice of loop/index order for all contraction path terms.
@@
A fully-fused loop nest and corresponding tree (forest) is obtained by fusion of the path graphs (loop nests) corresponding to each term in $A$. In particular, all common leading indices in subsequent terms are fused. We introduce a peeling primitive for fully fused loop nests to formally define the tree structure. Peeling a fully fused loop nest removes the first outermost loop nest.
@@
Given loop order $A=(A_{1},\ldots,A_{N})$, choose $r\in\{1,\ldots,N\}$ to be the largest such that $A_{1}[1]=A_{2}[1]=\cdots=A_{r}[1]$. Peeling $A$ yields two loop orders $A^{(1)}=(A_{1}[2:\ ],\ldots,A_{r}[2:\ ])$ and $A^{(2)}=(A_{r+1},\ldots,A_{N})$ (where $X[2:\ ]$ denotes the subsequence of all elements in $X$ except the first element and is omitted if $X$ has size 1).
The loop nest tree or forest can then be constructed from the representation $A=(A_{1},\ldots,A_{N})$ by peeling $A$ iteratively.
@@
Given an index order $A=(A_{1},\ldots,A_{n})$, the corresponding fully-fused loop nest forest $\mathcal{F}(A)=(V,E)$ is constructed as follows. Initialize $V$ as one vertex corresponding to loop index $A_{1}[1]$, then apply peeling iteratively. At each peeling step, add vertices to $V$ for $A^{(1)}$ and $A^{(2)}$ (unless they are zero-sized) connecting $A^{(1)}$ to the vertex representing $A$ and $A^{(2)}$ to its parent (if any).
With this definition, we now also define the effect of the peeling operation on a loop nest tree.
@@
Given a loop nest index order $A$ for contraction path $(T,L)$ and the corresponding fully-fused loop nest tree $\mathcal{F}(A)=(V,E)$, peeling removes the root vertex (index $r$) of the tree. If the root has $k$ children, the resulting independent subtrees are associated with index orders $B^{(1)},\ldots,B^{(k)}$, each of which computes a contraction path for distinct subsets of terms $L^{(1)},\ldots,L^{(k)}\subseteq\hat{L}$, where $\hat{L}$ is defined by removing the index $r$ from all index sets in $L$. The contraction path tree for the $i$th index order, $T^{(i)}$, is given by removing all vertices from $T$ except those corresponding to terms computed in $L^{(i)}$ and their children (inputs).
@@ 4.2.2. General Cost Function
In general, the execution time of a particular fully-fused loop nest tree may depend on architecture or data sparsity in ways that are impractical to fully model and require enumeration and execution. On the other hand, for a simple cost function, e.g., computational cost [^2] Since the same contraction path is being considered, all fully-fused loop nest trees have the same asymptotic complexity in tensor size, but order and fusion have an affect on lower-order cost terms. or intermediate buffer size, the search space can be explored more systematically and efficiently. However, more sophisticated cost functions, which take into account metrics such as cache-efficiency or parallelizability are also of clear interest. We now define a class of functions which we can optimize efficiently, requiring separability of cost according to the structure of the loop nest tree.
@@ 4.2.3. Maximum Buffer Size
We now provide a tree-separable cost function to compute the maximum dimension of the intermediate tensors/buffers produced in the execution of a fully fused loop nest.
@@
Consider a fully fused loop nest tree $\mathcal{F}(A)$ for loop order $A$ with contraction path $(T,L)$, where $T=(V,E)$. Let $B^{(1)},\ldots,B^{(k)}$ be the loop nest orders for subtrees obtained after peeling $\mathcal{F}(A)$ and $(T^{(i)},L^{(i)})$ be the corresponding contraction path for each $B^{(i)}$. Let $Z\subseteq E$ be the set of edges in the contraction path (oriented towards the root) connecting a node that corresponds to a term $L_{u}\in B^{(i)}$ to another, $L_{v}\in B^{(j)}$ with $i\neq j$. The maximum buffer dimension used in the fully fused loop nest is given by $f_{\varphi,\max}(T,L,A)$ where $f_{\varphi,\max}$ is a tree-separable cost function defined with $\varphi_{T,L,r}(x)=\max(\rho(T,L,r),x)$,
where $\rho(T,L,r)=\max_{(L_{u},L_{v})\in Z,L_{u}=(K_{1},K_{2},K_{3})}|K_{3}|$.
@@ 4.2.4. Total Number of Cache Misses
Algorithm 1 Algorithm to find cost-optimal index order for terms in a given contraction path
@@
3:Output: Two index orders, $A$ and $B$, for $(T,L)$, so that $A$ has minimal cost ($f_{\varphi,\oplus}(T,L,A)$) among all index orders for $(T,L)$ and $B$ has minimal cost among all index orders for $(T,L)$ that yield a loop nest tree $\mathcal{F}(B)$ with a different root than $\mathcal{F}(A)$.
@@
23:                  $\delta\leftarrow\varphi_{T,L,q}\Big{(}f_{\varphi,\oplus}(T^{(X)},L^{(X)},A^{(X)})\Big{)}\oplus f_{\varphi,\oplus}(T^{(Y)},L^{(Y)},A^{(Y)})$LCommentUpdate lowest cost index orders
@@
28:                  $\delta_{B}\leftarrow\delta_{A}$; $B\leftarrow A$
29:                  $\delta_{A}\leftarrow\delta_{C}$; $A\leftarrow C$
30:            else if $\delta_{C}<\delta_{B}$ then
31:                  $\delta_{B}\leftarrow\delta_{C}$; $B\leftarrow C$
32:      return $(A,B)$
4.2.5. Algorithm to find a cost-optimal index order
@@
We give a proof of correctness and show how the subproblems of Algorithm 1 can be memoized to reduce its complexity. For both, it is helpful to enumerate the subproblems (calls to function ORDER) in terms of
@@
We now use induction on the size of these subproblems to prove correctness.
@@
Consider a contraction path $(T,L)$ and a tree-separable cost function $f$ specified by $\varphi_{T,L}$ and $\oplus$. ORDER($T,L,\varphi_{T,L,r}$) (Algorithm 1) returns two index orders, $A$ and $B$, for $(T,L)$, so that $A$ has minimal cost ($f_{\varphi,\oplus}(T,L,A)$) among all index orders for $(T,L)$ and $B$ has minimal cost among all index orders for $(T,L)$ that yield a loop nest tree $\mathcal{F}(B)$ with a different root than $\mathcal{F}(A)$.
@@
Since $(T^{(1)},L^{(1)})$ and $(T^{(2)},L^{(2)})$ are contained and smaller (as defined in our inductive hypothesis) than $(T,L)$, Algorithm 1, when considering root vertex $r$, would return the minimal cost index order for both subproblems. Further, the cost of $A^{\prime}$ would be computed correctly on line 23 of the Algorithm. Since the algorithm instead found $A$ to have a lower cost, we have derived a contradiction. Given optimality of $A$, its trivial to check that the given optimality condition for $B$ is maintained. ∎
We now consider the execution cost of Algorithm 1, with the cost of each subproblem memoized. For $N$ ordered terms and $m$ total indices, there are $O((m!)^{N})$ index orders (loop nests). Algorithm 1 needs to consider all subsequences of the $N$ terms and all subsets of the $m$ indices, yielding $O(N^{2}2^{m})$ subproblems. Each subproblem considers all choices of root index and prefixes of terms that contain that index to iterate over. Thus the cost per subproblem is $O(mN)$ and the overall complexity of the algorithm is $O(N^{3}2^{m}m)$.
@@ 5. SpTTN-Cyclops Framework
We build a runtime framework for SpTTN kernels, which searches for cost-optimal loop nests using the methodology/algorithm introduced in Section 4 and executes the resulting loop nests. Specifically, the framework considers all contraction paths with optimal asymptotic complexity (max loop nest depth). For each contraction path, we restrict loop orders to those in which the indices of the sparse tensor are iterated over in the order in which they are stored in the CSF tree. We select the minimum cost loop nest among these using Algorithm 1. While the framework may use different cost functions and employ autotuning, in the experiments, we use a tree-decomposable cost metric that selects the loop nest with the maximum number of independent dense loops with bounded buffer dimension. This choice is made to use BLAS kernels as much as possible while maintaining a bounded amount of storage.

### v2

@@ 3.1. Loop Nests and Loop Nest Forests
For a contraction of $N+1$ tensors, a contraction path is given by a depth-first postordering of a $12N+$ 1-node binary tree $T$ where the $N+1$ leaves are the input tensors, and each internal node corresponds to the contraction of a pair of input tensors and/or intermediates, so all non-leaf nodes have exactly two children. We represent a contraction path by the tree $T$ and an ordered collection of index set 3-tuples, $L=(L_{1},\dots,L_{N})$, where each $L_{i}$ contains the indices of the tensor operands and output at each of the $N$ internal tree nodes.
Note that while a contraction path is defined above based on a tree, this tree is different tree from a loop nest tree. In a loop nest tree, each node corresponds to a loop and each leaf is a term in the contraction path. Hence, a node in the contraction path tree corresponds to a leaf in the loop nest tree. Figure 2(a) shows a contraction path tree for an order 4 TTMc kernel.
Figure 2. An order 4 TTMc kernel $\mathcal{S}_{irst}=\mathcal{T}_{ijkl}\cdot\mathcal{U}_{jr}\cdot\mathcal{V}_{ks}\cdot\mathcal{W}_{lt}$, where (a) represents the contraction path tree ($T$) with $L=((ijkl,lt,ijkt),(ijkt,ks,ijst),(ijst,jr,irst))$, and (b) shows the path graphs corresponding to the contraction path terms, fused to obtain a fully fused loop nest tree.
In a valid loop nest forest, all indices in a contraction term should be loop indices on the path between the corresponding leaf and the tree root, and the path should contain no additional or repeated indices. We refer to this order of loop indices as the loop order of the contraction term. For example, in Figure 2(b), the loop order of the first term, $\mathcal{T}_{ijkl}\cdot\mathcal{W}_{lt}\rightarrow\mathcal{X}_{ijkt}$, is $(i,j,k,l,t)$. If a vertex has multiple leaves in its subtree, the loop associated with that vertex contains all the contraction terms in that subtree.
@@
We say a loop nest tree is fully-fused if no vertex has two consecutive children that correspond to the same index. A fully-fused loop nest and the corresponding tree is obtained by fusion of the path graphs (loop nests) corresponding to each term in $A$. In Figure 2(b), the path graphs corresponding to the contraction path terms are fused to obtain a fully-fused loop nest tree. A loop nest forest is fully-fused if adding a dummy vertex and connecting it to all roots in the forest yields a fully-fused loop nest tree.
3.2. Intermediate Tensors
3.3. Contraction Path and Loop Order
The contraction path affects the asymptotic complexity (loop depth) and memory requirements (intermediate tensor sizes) of the computation. For example, consider the various ways to compute the TTMc kernel as shown in Figure 1. In one of the chosen contractions paths, tensors $\mathcal{T}$ and $\mathcal{V}$ are contracted first and the result is then contracted with $\mathcal{U}$. The computation has a maximum loop depth of 4 (Figures 1(a), 1(b) and 1(c)). A different contraction path of the same kernel, where tensors $\mathcal{U}$ and $\mathcal{V}$ are contracted first and then the result is contracted with $\mathcal{T}$, yields a maximum loop depth of 5 (Figure 1(d)).
Similarly, for a given contraction path, the ordering of vertices in the path graphs before fusing them, affects the intermediate tensor sizes and other cost metrics of interest. In the previous example of the TTMc kernel, consider the iteration graph in Figure 1(a) and its fully-fused variant in Figure 1(b). Indices $i$, $j$ and $s$ are common across the two trees in the iteration graph. We are able to fuse vertices $i$ and $j$ but not $s$ (loop order in the first path graph is $(i,j,k,s)$). This results in an intermediate tensor of size $S$ in Figure 1(b) (see Listing LABEL:lst:ttmc_fusion). But if the loop order in the first path graph is $(i,j,s,k)$, we can fuse vertices $i$, $j$ and $s$ in the iteration graph and obtain a fully-fused loop nest tree with an intermediate tensor of size 1 (scalar) (see Figure 1(c) and its corresponding loop nest in Listing LABEL:lst:ttmc_loop_order). In the next section, we seek to find cost-optimal loop nests for a given SpTTN kernel, where the cost is defined by a cost model, for example, the intermediate tensor size.
4. Finding Optimal SpTTN Kernels
To determine an efficient loop nest for an SpTTN kernel, we first present an approach to enumerate fully-fused trees and later present efficient algorithms to find an optimal tree for simple cost metrics.
4.1. Enumeration of Loop Nests
We seek to find cost-optimal loop nests for a given SpTTN kernel by enumerating only fully-fused loop nest forests and restrict our attention to dense multidimensional buffers (intermediate tensors). We decouple the enumeration into two steps: (1) enumeration of valid contraction paths for a given set of tensors and (2) enumeration of loop orders in the path graphs for a given contraction path.
4.1.1. Enumeration of Contraction Paths
Let the number of input tensors in the SpTTN be $n$. To enumerate contraction paths, we employ a function to pick and contract all combinations of two tensors from the list of input tensors. We then recurse over a new list constructed by replacing each pair of contracted tensors with the contraction output. This approach has been studied in the context of finding an optimal contraction path for dense tensor networks (Pfeifer et al., 2014). The cost can be analyzed by the recurrence relation, $T(n)={n\choose 2}\cdot T(n-1)$ and $T(2)=1$ (when there are two tensors to contract). The number of valid contraction paths for $n$ tensors is $O\left(\frac{(n!)^{2}}{n\cdot 2^{n}}\right)$.
In (Ibrahim et al., 2022), dynamic programming is used to find the cost-optimal contraction path (tree) given a fixed order of dense tensors to be contracted. This approach is analogous and complementary to our work of finding a cost-optimal loop nest tree for a given contraction path, which we present in Section 4.2.
4.1.2. Enumeration of Loop Orders for a Given Contraction Path
For a given contraction path, we construct a path graph for each term by picking a loop order for that term. The path graphs are then fused to obtain a fully fused loop nest tree. Each choice of loop order yields a different fused loop nest.
Let the set of indices in the $i$th term be $I_{i}$. The set of indices in the SpTTN is given by $I=\bigcup_{i=1}^{n-1}I_{i}$. We do an exhaustive search by enumerating all loop orders independently for each path graph and then considering all possible combinations of these orders. Since we do not allow any repeated indices in our path graphs, the loop nests generated in such an enumeration are unique and span all the possible loop nests for a given contraction path. The cardinality of this exhaustive search is given by $\prod_{i=1}^{n-1}|I_{i}|!$. We later restrict the search to only those loop orders that are consistent with the order of the indices of the sparse tensor, so if a term involves $k$ sparse indices, the number of possible orders for the term is only $|I_{i}|!/k!$.
4.1.3. Upper Bound on Loop Nests
For a given SpTTN, the number of loop nests we enumerate has an upper bound given by the product of the number of contraction paths and the number of loop orders for a given contraction path, i.e., $O\Big{(}\frac{(n!)^{2}\cdot\prod_{i=1}^{n-1}|I_{i}|!}{n\cdot 2^{n}}\Big{)}$. In the following section, we present a dynamic programming algorithm to prune the search space for loop order enumeration.
4.2. Algorithm to Find Cost-Optimal Loop Nests
Enumeration enables autotuning, but for analytic metrics of performance such as buffer size, more efficient search schemes are possible. Different contraction paths yield different fully-fused loop nests, hence we focus our attention to enumeration and search of loop nests for a particular contraction path. In TCE (Lam et al., 2000; Gao et al., 2007), dynamic programming is used to find the cost-optimal loop nest for dense tensor contractions, with one of the cost metrics being the intermediate tensor size. Our efficient search algorithm also employs dynamic programming, after decoupling order of terms from the tree structure. Given a fixed contraction path order (or a subsequence of the terms, which defines a subproblem), we seek to find a loop nest tree that minimizes a chosen cost metric.
[…]
background
@@ 4.2. Algorithm to Find Cost-Optimal Loop Nests
We introduce a peeling primitive for fully fused loop nests to formally define the tree structure. Peeling a fully fused loop nest removes the first outermost loop nest. In a fully-fused loop nest, the outermost loop should iterate over an index that appears in the first contraction, and include within it all subsequent contractions in the contraction path order until one does not include the index.
@@
Given loop order $A=(A_{1},\ldots,A_{N})$, choose $r\in\{1,\ldots,N\}$ to be the largest such that $A_{1}[1]=A_{2}[1]=\cdots=A_{r}[1]$. Peeling $A$ yields two loop orders $A^{(1)}=(A_{1}[2:\ ],\ldots,A_{r}[2:\ ])$ and $A^{(2)}=(A_{r+1},\ldots,A_{N})$ (where $A_{x}[2:\ ]$ denotes the subsequence of all elements in $A_{x}$ except the first element and is omitted if $A_{x}$ has size 1).
The loop nest tree or forest can then be constructed from the representation $A=(A_{1},\ldots,A_{N})$ by peeling $A$ iteratively and adding vertices for the two resulting loop orders (if not empty).
@@
Given an loop order $A=(A_{1},\ldots,A_{n})$, the corresponding fully-fused loop nest forest $\mathcal{F}(A)=(V,E)$ is constructed as follows. Initialize $V$ as one vertex corresponding to loop index $A_{1}[1]$, then apply peeling iteratively. At each peeling step, add vertices to $V$ for $A^{(1)}$ and $A^{(2)}$ (unless they are zero-sized) connecting $A^{(1)}$ to the vertex representing $A$ and $A^{(2)}$ to its parent (if any).
To work with analyzing loop nest forests, it also helps to think about the effect of peeling the loop order on the loop nest tree associated with the loop order.
@@
Given a loop nest loop order $A$ for contraction path $(T,L)$ and the corresponding fully-fused loop nest tree $\mathcal{F}(A)=(V,E)$, peeling removes the root vertex (index $r$) of the tree. If the root has $k$ children, the resulting independent subtrees are associated with loop orders $B^{(1)},\ldots,B^{(k)}$, each of which computes a contraction path for distinct subsets of terms $L^{(1)},\ldots,L^{(k)}\subseteq\hat{L}$, where $\hat{L}$ is defined by removing the index $r$ from all index sets in $L$. The contraction path tree for the $i$th loop order, $T^{(i)}$, is given by removing all vertices from $T$ except those corresponding to terms computed in $L^{(i)}$ and their children (inputs).
@@ 4.2.1. General Cost Function
[…]
@@ 4.2.2. Maximum Buffer Size
We now provide a tree-separable cost function to compute the maximum dimension of the intermediate tensors/buffers produced in the execution of a fully fused loop nest. We interchangeably use the terms intermediate tensor and buffer.
@@
Consider a fully fused loop nest tree $\mathcal{F}(A)$ for loop order $A$ with contraction path $(T,L)$, where $T=(V,E)$. Let $B^{(1)},\ldots,B^{(k)}$ be the loop nest orders for subtrees obtained after peeling $\mathcal{F}(A)$ and $(T^{(i)},L^{(i)})$ be the corresponding contraction path for each $B^{(i)}$. Let $Z\subseteq E$ be the set of edges in the contraction path (oriented towards the root) connecting a node that corresponds to a term $L_{u}\in B^{(i)}$ to another, $L_{v}\in B^{(j)}$ with $i\neq j$. The maximum buffer dimension used in the fully fused loop nest is given by $f_{\varphi,\max}(T,L,A)$ where $f_{\varphi,\max}$ is a tree-separable cost function defined as $\varphi_{T,L,r}(x)=\max(\rho(T,L,r),x)$, with $\rho(T,L,r)=\max_{(L_{u},L_{v})\in Z,L_{u}=(K_{1},K_{2},K_{3})}|K_{3}|$.
@@ 4.2.3. Total Number of Cache Misses
Algorithm 1 Algorithm to find cost-optimal loop order for terms in a given contraction path
@@
3:Output: Two loop orders, $A$ and $B$, for $(T,L)$, so that $A$ has minimal cost ($f_{\varphi,\oplus}(T,L,A)$) among all loop orders for $(T,L)$ and $B$ has minimal cost among all loop orders for $(T,L)$ that yield a loop nest tree $\mathcal{F}(B)$ with a different root than $\mathcal{F}(A)$.
@@
23:             $\delta\leftarrow\varphi_{T,L,q}\Big{(}f_{\varphi,\oplus}(T^{(X)},L^{(X)},A^{(X)})\Big{)}\oplus f_{\varphi,\oplus}(T^{(Y)},L^{(Y)},A^{(Y)})$LCommentUpdate lowest cost loop orders
@@
28:             $\delta_{B}\leftarrow\delta_{A}$; $B\leftarrow A$; $\delta_{A}\leftarrow\delta_{C}$; $A\leftarrow C$
29:         else if $\delta_{C}<\delta_{B}$ then
30:             $\delta_{B}\leftarrow\delta_{C}$; $B\leftarrow C$
31:    return $(A,B)$
@@
We now provide a proof of correctness and show how the subproblems of Algorithm 1 can be memoized to reduce its complexity. For both, it is helpful to enumerate the subproblems (calls to function ORDER) in terms of
@@
We use induction on the size of these subproblems to prove correctness.
@@
Consider a contraction path $(T,L)$ and a tree-separable cost function $f$ specified by $\varphi_{T,L}$ and $\oplus$. ORDER($T,L,\varphi_{T,L,r}$) (Algorithm 1) returns two loop orders, $A$ and $B$, for $(T,L)$, so that $A$ has minimal cost ($f_{\varphi,\oplus}(T,L,A)$) among all loop orders for $(T,L)$ and $B$ has minimal cost among all loop orders for $(T,L)$ that yield a loop nest tree $\mathcal{F}(B)$ with a different root than $\mathcal{F}(A)$.
@@
Since $(T^{(1)},L^{(1)})$ and $(T^{(2)},L^{(2)})$ are contained and smaller (as defined in our inductive hypothesis) than $(T,L)$, Algorithm 1, when considering root vertex $r$, would return the minimal cost loop order for both subproblems. Further, the cost of $A^{\prime}$ would be computed correctly on line 23 of the Algorithm. Since the algorithm instead found $A$ to have a lower cost, we have derived a contradiction. Given optimality of $A$, its trivial to check that the given optimality condition for $B$ is maintained. ∎
We now consider the execution cost of Algorithm 1, with the cost of each subproblem memoized. For $N$ ordered terms and $m$ total indices, there are $O((m!)^{N})$ loop orders (loop nests). Algorithm 1 needs to consider all subsequences of the $N$ terms and all subsets of the $m$ indices, yielding $O(N^{2}2^{m})$ subproblems. Each subproblem considers all choices of root index and prefixes of terms that contain that index to iterate over. Thus the cost per subproblem is $O(mN)$ and the overall complexity of the algorithm is $O(N^{3}2^{m}m)$.
[…]
algorithm
@@ 5. SpTTN-Cyclops Framework
We build a runtime framework for SpTTN kernels, which searches for cost-optimal loop nests using the methodology/algorithm introduced in Section 4 and executes the resulting loop nests. Specifically, the framework first considers all contraction paths with optimal asymptotic complexity. For each contraction path, we restrict loop orders to those in which the indices of the sparse tensor are iterated over in the order in which they are stored in the CSF tree. We select the minimum cost loop nest among these using Algorithm 1. If the framework cannot find a loop nest that fits within the constraints set by the cost model, it iterates over the contraction paths with suboptimal asymptotic complexity until it finds a loop nest that adheres to the constraints. While the framework may use different cost functions and employ autotuning, in the experiments, we use a tree-decomposable cost metric that selects the loop nest with the maximum number of independent dense loops with bounded buffer dimension. This choice is made to use BLAS kernels as much as possible while maintaining a bounded amount of storage.
1T_csf = CSF(${\mathcal{T}\_ijkl}$)
2for (i,T_i) in T_csf:
3 for (j,T_ij) in T_i:
4 $\mathcal{Y}$ = 0 // reset intermediate tensor
5 for (k,T_ijk) in T_ij:
6 $\mathcal{X}$ = 0 // reset intermediate tensor
7 for l in T_ijk:
8 xAXPY(T, t_ijkl, $\mathcal{W}$[l,:], 1, $\mathcal{X}$, 1)
9 xGER(T, S, 1, $\mathcal{X}$, 1, $\mathcal{V}$[k,:], 1, $\mathcal{Y}$, T)}
10 for r in range(R): // dense loop
11 xAXPY((S * T), $\mathcal{Y}$, $\mathcal{U}$[j,r], 1, $\mathcal{S}$[i,r,:,:], 1)
Figure 3. Loop nest for an order 4 TTMc kernel. Loop $r$ of contraction 3 is not via recursion but is generated as a loop by metaprogramming. Contractions 1 and 3 are offloaded to BLAS-1, and contraction 2 is offloaded to a BLAS-2 kernel.
5.1. Algorithm to Generate and Execute Loop Nests
Given a fully fused loop nest tree, in Algorithm 2 we present a runtime algorithm to generate loop nests and execute the contractions. We represent the tree with a sequence of terms (leaves) and a list per term representing the loop order (vertices). This representation is sufficient for the algorithm to infer the structure of a fully fused loop nest tree. We use Algorithm 2 in two stages. In the first stage, we preprocess the fully fused loop nest tree and add hooks to (1) generate nested loops for the dense indices using metaprogramming, (2) identify independent dense loops that can be offloaded to BLAS like kernels. We also allocate memory for the intermediate tensors in this stage. In the second stage, we compute the kernel by executing the preprocessed fully fused loop nest tree. We check for hooks in Line 2 and offload the computation accordingly.